# RoToR: Towards More Reliable Responses for Order-Invariant Inputs

### v2

@@ 4.1 Baselines
Original causal LM with no modifications (Orig.) were compared, which processes text sequentially. Also, we compare RoToR against previous zero-shot order-invariant LMs discussed in Sec. 3.1, namely PCW Ratner et al. (2023), PINE Wang et al. (2024), and Set-Based Prompting McIlroy-Young et al. (2024). We use the LLaMA 3.1 AI (2024) 8B-Instruct444meta-llama/Meta-Llama-3.1-8B-Instruct, Qwen1.5-4B-Chat and Qwen1.5-7B-Chat [^5] Qwen/Qwen1.5-4/7B-Chat as our backbone model for all of our experiments. As our method doesn’t need training, a single A6000 GPU was sufficient to run all of the experiments except for the Llama-3.1-70B-Instruct model (Appendix, Tab. 6). We also conduct experiments on a subset of benchmarks (LitM and MMLU) on the runtime latency, perplexity, and collision rate of PINE and RoToR, to further validate our claims on Sec. 3.2.
@@ 4.2 Benchmarks with listwise inputs
We experiment with three benchmarks involving real-world listwise input data. Examples of exact inputs and outputs are provided in Appendix A.7. All reported scores are rounded to the nearest tenth, except for the standard deviation (rounded to the second decimal place).
Knowledge Graph QA (KGQA) In KGQA tasks, the model takes facts over knowledge graphs represented as (subject, relation, object), and answers the given question based on the given facts. We basically follow the KGQA dataset preprocessing and evaluation setup from  Baek et al. (2023), which uses Mintaka Sen et al. (2022) with Wikidata for knowledge source, and use the Exact Match (EM), Accuracy (Acc), and F1 score metric for evaluation. We also use MPNet Song et al. (2020) as a dense retriever to retrieve top-k facts over each question, and experiment with segment size of 30 and 50. Replication details and example dataset format are at Appendix Sec. A.3 and Fig. 10. Along with measuring the performance of the initial input ordering, we report performance after we shuffle the order of segments with 3 different seeds to see shuffle robustness.
Lost in the middle (LitM) We use the Lost in the Middle (LitM) benchmark Liu et al. (2024), which draws from 2655 queries in the Natural Questions (NQ) dataset. It provides sets of (10, 20, 30) passages, placing the gold passage at predetermined positions (e.g., 0, 4, 9) and filling the remaining slots with irrelevant passages. Following  Liu et al. (2024), the best_subspan_em metric is used. Experiments on LitM found that eliminating the effect of index bias is another important detail for measuring true order robustness: (Appendix Sec. B). Thus, we report experiments with index bias eliminated. The exact prompt and full results including index bias is reported at Appendix Fig. 8 and Sec. A.1.
@@ 5 Results & Analysis
We observe that shuffling input segments leads to non-trivial performance degradations in the original model, which exhibits a statistically significant performance drop on our experimented dataset (two-tailed t-test, p < 0.05, Appendix C). In contrast, our proposed RoToR model does not show a statistically significant difference in performance before and after shuffling, indicating that it is more robust against such perturbations. On LitM (Tab. 1), we notice PCW and Set-Based Prompting has impractical performance, with PINE degrading heavily as number of documents ($k$) increases, while RoToR is less affected. On KGQA (Tab. 2), we show RoToR outperform PINE with lower standard deviation across shuffled segments, consistent with different model architectures.
@@
Runtime latency (s)
Perple-
xity ($\downarrow$)
Collision
Rate($\downarrow$)
MMLU LitM LitM
data count 14015 2655 2655
# segments 4 10 30 20 30
PINE 7,371 18,551 41,664 6.91 42.3%
RoToR 6,608 14,264 23,569 6.65 0 (None)
Reduced % 10.4% 23.1% 43.4%
Table 4: Compared to PINE, RoToR with lexical sort is faster, has lower perplexity, and have zero collision rate between segments. We use a subset of LitM with index bias removed and gold index = 0.
Experiments against comparing RoToR with PINE (Tab. 4) analyze the following: Runtime, scalability: Actual inference times (Appendix Sec. A.6) find that RoToR outperforms PINE substantially, with efficiency gains increasing alongside $n$. For instance, on LitM (30 docs), RoToR achieves a 43% reduction in total runtime. Practical scalability with increasing $k$ is critical, but we find that previous order-invariant LMs struggle handling larger $k$ (on KGQA and LitM). In contrast, RoToR shows better performance with improved efficiency and robustness. Perplexity: Lower generation perplexity indicates input representations are closer to in-distribution. On the same LitM dataset, RoToR’s reduced perplexity implies its positional ID assignment effectively mitigates out-of-distribution effects. Collision Rate: PINE’s similarity-based ordering often collides: on average, only 17.3 of 30 similarity values are unique, causing 42% of the segments to be indistinguishable and thus breaking invariance. In contrast, RoToR with lexical sorting only collides if the segment texts are identical. On LitM, this yields zero collisions, preserving full invariance.
@@
MMLU (Tab. 3) is a representative of a task that involves not only order-invariant, but also order-sensitive (e.g., "None of the above"), inputs. Therefore, single use of order-invariant models does not always outperform the original model, limiting applicability of order-invariant models to practical listwise tasks, i.e., we observe significant performance drops for Set-based Prompting in MMLU, falling short of half the performance of the original model on initial ordering. However, using RoToR with Selective Routing to handle order-sensitive inputs outperforms, or is at least competitive as the original model in all possible orderings of candidate choices. Selective Routing improves the generalizability and extends the applicability on practical listwise tasks by adaptively handling order-sensitive inputs. The RoToR + Selective Routing (Optimal) performance on Tab. 3 was evaluated using a relaxed accuracy metric based on the union of predictions from the original and the invariant (RoToR-lexical) model. This improves significantly, which highlights the potential of Selective Routing for further accuracy gains through optimizing design choices on routing methods, which we plan to explore in future work.
@@ 6 Conclusion
Our work addresses order-invariance in listwise inputs by identifying core issues in distribution mismatch and adaptive handling of mixed inputs. Our proposed RoToR provides a stable zero-shot order-invariant solution that reduces the complexity of positional ID modification, while Selective Routing adaptively routes between invariant and sensitive LMs to handle real-world scenarios. Together, these methods demonstrate improved performance and reliability on LitM, KGQA, and MMLU benchmarks.
@@ 7 Limitations
Appendix A Appendix
A.1 Full results on the Lost in the Middle Benchmark
@@
ndoc = 10 20
gold idx = 0 4 9 avg. 0 4 9 14 19 avg.
Orig. 66.2 65.7 65.7 65.9 65.2 64.3 65.0 66.2 64.8 65.2
PINE 67.9 67.8 67.5 67.7 65.9 65.7 65.9 65.8 65.5 66.1
RoToR-reranking 68.9 69.0 68.8 68.9 67.5 67.5 67.7 67.5 67.6 67.8
RoToR-lexical 69.6 69.5 69.3 69.5 67.6 67.8 67.8 67.7 67.9 68.0
Table 6: Reporting the performance with the Llama-3.1-70B-Instruct model on a subset of lost-in-the middle benchmark (without index bias) performance. We see that the gains are consistant with the models we mainly investigated (Llama-3.1-8B-Instruct)
@@
Results on other models.
Table  6 shows results in different model (Llama-3.1-70B-Instruct). We can see that the trend continues for the 70B model, implying the robustness of RoToR.
A.2 Illustration of the global sorting method
@@
We show the two different global sorting algorithms presented in our paper at Fig. 6.
A.3 Details about preprocessing and evaluation of datasets
A.3.1 General.
@@ A.3 Details about preprocessing and evaluation of datasets
A.3.2 Knowledge Graph Question Answering
@@
A.3.3 Lost in the Middle
Specifically, we use the dataset provided in the official repository666github.com/nelson-liu/lost-in-the-middle, use the same prompt as the llama 2 chat model with only the instruction tokens adjusted to llama 3 (removed [Inst] and changed to <|begin_of_text|> and etc.,), and evaluate using the best_supspan_em metric.
A.3.4 MMLU
@@
A.4 Further impact scenarios on general conversation.
@@
A.5 Selection of $\alpha$ for Selective Routing on MMLU
@@
A.6 Replication details on the runtime experiment
@@
A.7 Input data examples
@@
Appendix B Why is removing index bias an important detail for invariant models to be effective?
@@
Appendix C Statisticial significance before and after shuffling segments

### v3

@@ 4.1 Baselines
Original causal LM with no modifications (Orig.) were compared, which processes text sequentially. Also, we compare RoToR against previous zero-shot order-invariant LMs discussed in Sec. 3.1, namely PCW Ratner et al. (2023), PINE Wang et al. (2024), and Set-Based Prompting McIlroy-Young et al. (2024). We use the LLaMA 3.1 AI (2024) 8B-Instruct666meta-llama/Meta-Llama-3.1-8B-Instruct 70B-Instruct, Qwen1.5-4B-Chat, Qwen1.5-7B-Chat [^7] Qwen/Qwen1.5-4/7B-Chat, and Qwen1.5-72B-Chat as our backbone model for our experiments. As our method doesn’t need training, a single A6000 GPU was sufficient to run all of the experiments except for the Llama-3.1-70B-Instruct and Qwen1.5-72B-Chat model. We also conduct experiments on a subset of benchmarks (LitM and MMLU) on the runtime latency, perplexity, and collision rate of PINE and RoToR, to further validate our claims on Sec. 3.2.
@@ 4.2 Benchmarks with listwise inputs
We experiment with three benchmarks involving real-world listwise input data. Examples of exact inputs and outputs are provided in Appendix G. All reported scores are rounded to the nearest tenth, except for the standard deviation (rounded to the second decimal place).
Knowledge Graph QA (KGQA) In KGQA tasks, the model takes facts over knowledge graphs represented as (subject, relation, object), and answers the given question based on the given facts. We basically follow the KGQA dataset preprocessing and evaluation setup from  Baek et al. (2023), which uses Mintaka Sen et al. (2022) with Wikidata for knowledge source, and use the Exact Match (EM), Accuracy (Acc), and F1 score metric for evaluation. We also use MPNet Song et al. (2020) as a dense retriever to retrieve top-k facts over each question, and experiment with segment size of 30 and 50. Replication details and example dataset format are at Appendix Sec. C and Fig. 10. Along with measuring the performance of the initial input ordering, we report performance after we shuffle the order of segments with 3 different seeds to see shuffle robustness.
Lost in the middle (LitM) We use the Lost in the Middle (LitM) benchmark Liu et al. (2024), which draws from 2655 queries in the Natural Questions (NQ) dataset. It provides sets of (10, 20, 30) passages, placing the gold passage at predetermined positions (e.g., 0, 4, 9) and filling the remaining slots with irrelevant passages. Following  Liu et al. (2024), the best_subspan_em metric is used. Experiments on LitM found that eliminating the effect of index bias is another important detail for measuring true order robustness: (Appendix Sec. H). Thus, we report experiments with index bias eliminated. The exact prompt and full results including index bias is reported at Appendix Fig. 8 and Sec. A.
@@ 5 Results & Analysis
We observe that shuffling input segments leads to non-trivial performance degradations in the original model, which exhibits a statistically significant performance drop on our experimented dataset (two-tailed t-test, p < 0.05, Appendix I). In contrast, our proposed RoToR model does not show a statistically significant difference in performance before and after shuffling, indicating that it is more robust against such perturbations. On LitM (Tab. 1), we notice PCW and Set-Based Prompting has impractical performance, with PINE degrading heavily as number of documents ($k$) increases, while RoToR is less affected. On KGQA (Tab. 2), we show RoToR outperform PINE with lower standard deviation across shuffled segments, consistent with different model architectures.
@@
Model Benchmark PINE RoToR Reduction
(a) Overhead FLOPs, relative to original model
Llama-3.1- 8B-Instruct MMLU, $N=4$ 0.59$\times$ 0.55$\times$ 7.6%
LitM, $N=10$ 7.07$\times$ 4.81$\times$ 31.9%
LitM, $N=30$ 22.43$\times$ 15.05$\times$ 32.9%
Llama-3.1- 70B-Instruct KGQA, $N=30$ 1.27$\times$ 0.94$\times$ 26.0%
KGQA, $N=50$ 1.82$\times$ 1.29$\times$ 29.0%
Qwen1.5- 72B-Chat KGQA, $N=30$ 0.45$\times$ 0.01$\times$ 98.0%
KGQA, $N=50$ 0.58$\times$ 0.03$\times$ 94.8%
(b) End-to-end latency (s)
Llama-3.1- 70B-Instruct LitM, $N=10$ 57,352 44,219 22.9%
LitM, $N=20$ 87,091 58,680 32.6%
Llama-3.1- 8B-Instruct MMLU, $N=4$ 7,371 6,608 10.4%
LitM, $N=10$ 18,551 14,264 23.1%
LitM, $N=30$ 41,664 23,569 43.4%
(c) Perplexity & Collision rate, (on LitM)
Llama-3.1- 8B-Instruct Perplexity ($N=20$) 6.91 6.65
Collision rate ($N=30$) 42.3% 0 (None)
Table 4: Unified efficiency comparison of RoToR vs. PINE, reporting (a) Additional FLOPs, (b) Latency, and (c) Perplexity & Collision rate. Columns list each metric for PINE and RoToR, and the relative reduction. Yellow rows separate sub-sections.
Experiments against comparing RoToR with PINE (Tab. 4) we analyze FLOPs:  [^8] We used the FlopsProfiler of the DeepSpeed library to measure FLOPs. RoToR consistently reduces the floating point operations overhead across segment counts and different model backbones compared to PINE. This is because RoToR does not require computing additional attention scores: it only performs tensor operations for circular arrangement. In contrast, PINE requires more cost due to attention-based reassignment. Runtime, scalability: Actual inference times (Appendix Sec. F) find that RoToR outperforms PINE substantially, with efficiency gains increasing alongside $n$. For instance, on LitM (30 docs), RoToR achieves a 43% reduction in total runtime. Practical scalability with increasing $k$ is critical, but we find that previous order-invariant LMs struggle handling larger $k$ (on KGQA and LitM). In contrast, RoToR shows better performance with improved efficiency and robustness. Perplexity: Lower generation perplexity indicates input representations are closer to in-distribution. On the same LitM dataset, RoToR’s reduced perplexity implies its positional ID assignment effectively mitigates out-of-distribution effects. Collision Rate: PINE’s similarity-based ordering often collides: on average, only 17.3 of 30 similarity values are unique, causing 42% of the segments to be indistinguishable and thus breaking invariance. In contrast, RoToR with lexical sorting only collides if the segment texts are identical. On LitM, this yields zero collisions, preserving full invariance.
@@
MMLU (Tab. 3) is a representative of a task that involves not only order-invariant, but also order-sensitive (e.g., "None of the above"), inputs. Therefore, single use of order-invariant models does not always outperform the original model, limiting applicability of order-invariant models to practical listwise tasks, i.e., we observe significant performance drops for Set-based Prompting in MMLU, falling short of half the performance of the original model on initial ordering. However, using RoToR with Selective Routing to handle order-sensitive inputs outperforms, or is at least competitive as the original model in all possible orderings of candidate choices. Selective Routing improves the generalizability and extends the applicability on practical listwise tasks by adaptively handling order-sensitive inputs. The RoToR + Selective Routing (Oracle) performance on Tab. 3 was evaluated using a relaxed accuracy metric based on the union of predictions from the original and the invariant (RoToR-lexical) model. This improves significantly, which highlights the potential of Selective Routing for further accuracy gains through optimizing design choices on routing methods, which we plan to explore in future work. Additional analysis on the selection ratio of Selective Routing is reported at Appendix Section L.
@@
Extension to LLMs, other scenarios
Experiments on Llama-3.1-70B-Instruct and Qwen-1.5-72B-Chat for LitM and KGQA show consistent and significant improvements over both the original implementation and the PINE baseline, demonstrating RoToR’s generalizability to larger-scale LLMs. We further evaluate robustness on longer-context inputs (LongBench-2WikiMultiHopQA Bai et al. (2024) in Appendix Section J) and different task templates (KGQA template swap in Appendix Section K). Results indicate that our method retains benefits across longer input scenarios and diverse task templates, including those without explicit input format requirements.
@@ 6 Conclusion
Our work addresses order-invariance in listwise inputs by identifying core issues in distribution mismatch and adaptive handling of mixed inputs. Our proposed RoToR by modifying self-attention by global sorting and circular arrangement provides a stable zero-shot order-invariant solution that reduces the complexity of positional ID modification, while Selective Routing adaptively routes between invariant and sensitive LMs to handle real-world scenarios. Together, these methods demonstrate improved performance and reliability on LitM, KGQA, and MMLU benchmarks.
@@ 7 Limitations
8 Acknowledgments
We thank the Channel Corporation for providing GPU resources to run the experiments, and their AI team for providing helpful feedback. We thank the members of LDILab and Jinwoo Kim for their constructive comments and the anonymous reviewers for their valuable suggestions. We are also grateful to our former intern, Yezun Chung (KAIST CS), for assisting with experiments on early versions of the Selective Routing approach.
This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. RS-2021-II212068, Artificial Intelligence Innovation Hub) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)]
Appendix
Appendix A Full results on the Lost in the Middle Benchmark
@@
Appendix B Illustration of the global sorting method
@@
We show the three different global sorting algorithms presented in our paper at Fig. 6.
Appendix C Details about preprocessing and evaluation of datasets
C.1 General.
@@ Appendix C Details about preprocessing and evaluation of datasets
C.2 Knowledge Graph Question Answering
@@
C.3 Lost in the Middle
Specifically, we use the dataset provided in the official repository999github.com/nelson-liu/lost-in-the-middle, use the same prompt as the llama 2 chat model with only the instruction tokens adjusted to llama 3 (removed [Inst] and changed to <|begin_of_text|> and etc.,), and evaluate using the best_supspan_em metric.
C.4 MMLU
@@
Appendix D Further impact scenarios on general conversation.
@@
Appendix E Selection of $\alpha$ for Selective Routing on MMLU
@@
Appendix F Replication details on the runtime experiment
@@
Appendix G Input data examples
@@
Appendix H Why is removing index bias an important detail for invariant models to be effective?
@@
Appendix I Statisticial significance before and after shuffling segments
@@ Conclusion.
Appendix J Application to Long-Context Inputs
Benchmark and protocol.
To test whether RoToR consistantly performs well to inputs with longer contexts, we extend our evaluation to the LongBench Bai et al. (2024)2WikiMultihopQA task, whose multi-document questions yield input lengths from 5 k to 15 k tokens. Because our current RoToR and PINE implementations do not yet support advanced parallelization (e.g., FlashAttention or SDPA), GPU memory becomes prohibitive beyond $\sim$10 k tokens, especially for PINE, whose memory footprint is larger than ours. We therefore truncate the context at 10 k tokens, which already exceeds the lengths used in our main experiments.
Following the official LongBench guidelines [^12] Maximum generation length 32 and qa_f1_score (LongBench-E) as the evaluation metric., we segment each long context into 512-token “listwise” inputs, enabling a direct comparison with our listwise reranking pipeline. To ensure robustness, we test three input-order perturbations:
  1. 1.
    Initial: original chunk order $[1,2,3,4,5]$;
  2. 2.
    Reversed: $[5,4,3,2,1]$;
  3. 3.
    Center-flipped: first and last halves swapped, [3,2,1,5,4].
Results.
Table 9 reports F1 scores for the Original model variant, RoToR, and PINE with Llama 3.1-8B-Instruct and Qwen 1.5-7B-Chat. Across every ordering and token-length band, RoToR retains a clear advantage, while PINE fails to run when the longest (8 k+) chunks are kept. These findings confirm that RoToR scales to substantially longer inputs across different input-order perturbations.
Llama 3.1-8B-Instruct Qwen 1.5-7B-Chat
Order Method 0–4k 4–8k 8k+ Total 0–4k 4–8k 8k+ Total
Count 25 131 144 300 23 121 156 300
Initial (e.g., 1,2,3,4,5) Orig. 48.3 56.8 34.0 45.1 65.6 47.9 26.7 38.2
PINE 51.0 47.6 70.2 45.1
RoToR 59.0 52.7 41.8 48.0 75.7 47.8 31.0 41.2
Reversed (e.g., 5,4,3,2,1) Orig. 57.0 51.5 39.0 46.0 53.4 43.3 34.2 39.3
PINE 43.0 49.8 64.1 48.9
RoToR 59.0 52.0 41.0 47.3 72.8 47.6 30.8 40.8
Center flip (e.g., 3,2,1,5,4) Orig. 47.0 47.7 35.6 41.8 61.0 40.6 32.7 38.1
PINE 46.3 49.2 70.2 43.5
RoToR 59.0 52.5 41.5 47.8 77.1 47.3 30.9 41.0
Table 9: F1 scores (%) on LongBench–2WikiMultihopQA with $\sim$10k-token contexts. “Count” is the number of examples per length bucket; “–” denotes out-of-memory.
Take-aways.
Even without specialised long-context kernels, RoToR consistently outperforms both the original reranker and PINE, and remains robust to severe input-order perturbations. This suggests our approach can generalize to substantially longer inputs once memory and kernel constraints are alleviated.
Appendix K Robustness to task templates
Llama-3.1-8B-Instruct Qwen-1.5-4B-Chat Qwen-1.5-7B-Chat
$N=30$ $N=50$ $N=30$ $N=50$ $N=30$ $N=50$
Method Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1
Original template, Initial ordering
Template-swap
Original 50.0 44.1 51.8 50.2 44.3 51.9 31.1 27.8 34.9 31.7 28.3 35.3 31.4 27.6 35.2 32.0 28.0 35.7
PINE 52.0 45.7 52.9 52.0 45.3 52.8 31.9 28.8 35.8 31.7 28.6 35.4 31.9 28.5 36.0 31.5 28.2 35.7
RoToR 52.7 46.4 54.0 52.9 46.4 53.7 31.8 28.2 35.0 32.4 29.0 35.6 34.1 29.8 37.6 34.0 29.9 37.7
RoToR-MonoT5 51.5 45.2 52.6 52.5 45.7 53.1 32.4 29.0 36.3 32.6 29.3 35.9 32.9 28.5 36.4 32.6 28.5 36.3
RoToR-Freq. 52.3 46.2 53.7 52.9 46.5 53.8 31.9 28.4 35.4 32.4 28.8 35.6 34.0 29.9 37.4 33.8 29.6 37.4
Table 10: Results on the Mintaka (KGQA) dataset on different models, before (top block, also reported at main paper) and after (bottom block) the template-swap. N refers to number of top-k segments per query. RoToR variants consistently outperform the Original and PINE baselines, and their performance is stable under the swapped template, indicating robustness to instruction wording.
We evaluated whether the proposed selective-routing methods remain effective when the surrounding task instructions are re-phrased. Concretely, we performed a template-swap experiment on the KGQA benchmark, specifically on the initial ordering setup. The original prompt (Figure 10) began with
“Below are the facts in the form of triples meaningful to answer the question.”
and required the model to output only a JSON object. In the swapped template we replaced the first sentence with
“Below are knowledge statements expressed as triples meaningful to answer the question.”
leaving all other instructions unchanged. Table 10 reports the results. Across all three backbone models and both retrieval depths ($N{=}30,50$), RoToR and its variants retain similar absolute scores and continue to outperform both the Original and PINE baselines, indicating strong robustness to superficial wording changes in the task template.
Appendix L Additional Statistics on Selective Routing Assignment
Table 11 complements the main results in Table 3 by reporting the selection ratio, accounting for the percentage of evaluation queries for which the RoToR branch is chosen over the vanilla branch, under Selective Routing (SR). [^13] All figures are computed over the full evaluation set of 14,015 queries. We break the analysis down by (i) the global sorting strategy (Lexical, MonoT5, or Freq.), and (ii) three model backbones. The table distinguishes three order–based conditions:
Llama-3.1-8B-Instr. Qwen1.5-4B-Chat Qwen1.5-7B-Chat
Sorting Init. Rev. Avg. Init. Rev. Avg. Init. Rev. Avg.
Lexical 7.0 8.5 $7.3\!\pm\!0.8$ 5.9 6.2 $6.2\!\pm\!0.4$ 10.3 10.6 $9.9\!\pm\!0.6$
MonoT5 6.9 7.6 $6.7\!\pm\!1.5$ 8.0 12.5 $9.8\!\pm\!2.1$ 10.7 10.9 $10.7\!\pm\!0.7$
Freq. 6.4 6.7 $6.9\!\pm\!0.5$ 8.5 10.9 $9.4\!\pm\!1.6$ 10.7 11.1 $11.1\!\pm\!0.8$
Table 11: Selection ratio (%) of the RoToR variant under SR. Higher values indicate more frequent routing to RoToR.
For table 11, Init. refers to the original ordering (e.g., in abcd order), Rev. refers to the reversed ordering (e.g., in dcba but assigned as abcd), and Avg. is the average selection ratio for all possible (4!-1) re-orderings, with standard deviation. Empirically, the RoToR model tends to be selected more frequently under reversed orderings, whereas under the original ordering, the vanilla model is chosen slightly more often. The exact ratio varies by model and sorting strategy.